OpenStudy (anonymous):

Find the LCM w^3+10w^2+25w,w^2-10w

OpenStudy (anonymous):

w^3+10w^2+25w = w(w^2 + 10w + 25)= w (w+5)^2 w^2 - 10w = w(w - 10) LCM = w(w-10)(w+5)^2